# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ IV Cnoidal wave in a superfluid
As a particular application of the formalism developed in the previous section, we consider a 1D quasi-condensate Menotti and Stringari (2002) in a ring of length $L$ with no external potential, $V(x)=0$, so there is invariance under spatial translations. This type of model has been recently used to understand certain aspects of supersolidity due to its appealing simplicity Martone et al. (2021). We now look for solutions which simultaneously break both $U(1)$ and space-translational symmetry. In our notation, this is translated into $\boldsymbol{\alpha}=(\theta,x)$, $\mathbf{T}=(1,-i\partial_{x})$ and $\mathbf{Q}=(N,P)$, respectively. The Lagrange multipliers associated to the particle number $N$ and momentum $P$ are $\lambda^{\theta}=\mu$ and $\lambda^{x}=v$. The resulting GPG equation (23) then reads
@@
$H_{GP}\tilde{\Psi}_{0}=\left[\mu+\frac{v^{2}}{2}\right]\tilde{\Psi}_{0}\equiv\mu_{v}\tilde{\Psi}_{0}.$ (84)
@@
$\theta(x)$ $=$ $\int^{x}_{0}\frac{J}{n(x^{\prime})}\mathrm{d}x^{\prime}$
$=$ $\frac{J}{n_{1}\sqrt{n_{3}-n_{1}}}\Pi\left(\text{am}(\sqrt{n_{3}-n_{1}}x,\nu),m,\nu\right),$
@@
$\nu=\frac{n_{2}-n_{1}}{n_{3}-n_{1}},~{}m=1-\frac{n_{2}}{n_{1}},$ (86)
@@
$\mu_{v}=\frac{n_{1}+n_{2}+n_{3}}{2},~{}J=\sqrt{n_{1}n_{2}n_{3}}.$ (87)
@@
The density $n(x)$ of a cnoidal wave is a periodic function, with period
@@
Regarding the phase $\theta(x)$, since $n(x)$ is periodic, $J/n(x)$ is periodic. The integral of a periodic function is a periodic function modulo a constant slope $w_{\mathbf{n}}x$, given by its zero Fourier component. In this case,
@@
$L$ $=$ $\ell a=\frac{2K(\nu)\ell}{\sqrt{n_{3}-n_{1}}},~{}\ell\in\mathbb{N},$ (92)
$\frac{2\pi q}{L}$ $=$ $v+w_{\mathbf{n}},~{}q\in\mathbb{Z},$
$N$ $=$ $\int^{L}_{0}\mathrm{d}x~{}|\Psi_{0}(x)|^{2},$
$P$ $=$ $\int^{L}_{0}\mathrm{d}x~{}\Psi^{*}_{0}(-i\partial_{x})\Psi_{0},$
@@
$\frac{2\pi q}{L}$ $=$ $\frac{\bar{p}}{\bar{n}}+\sqrt{n_{1}n_{2}n_{3}}\left[\frac{\Pi(m,\nu)}{n_{1}K(\nu)}-\frac{1}{\bar{n}}\right],$
@@
$\bar{n}\equiv\frac{N}{L},~{}\bar{p}\equiv\dfrac{P}{L}.$ (94)
The procedure to numerically solve these equations is explained after Eq. (190). Once done, we can evaluate the energy density by
@@
and numerically compute the derivatives $\partial_{N}E=\partial_{\bar{n}}e$, $\partial_{P}E=\partial_{\bar{p}}e$, checking that indeed we recover $\partial_{N}E=\mu,~{}\partial_{P}E=v$. We note that the above formalism can be straightforwardly adapted to account for the presence of a flux within the ring.
@@
$z_{\theta}$ $=$ $\left[\begin{array}[]{r}-i\Psi_{0}\\ i\Psi_{0}^{*}\end{array}\right],~{}z_{x}=-\left[\begin{array}[]{l}\partial_{x}\Psi_{0}\\ \partial_{x}\Psi_{0}^{*}\end{array}\right],$ (100)
$z_{N}$ $=$ $\left[\begin{array}[]{l}\partial_{N}\Psi_{0}\\ \partial_{N}\Psi_{0}^{*}\end{array}\right],\,z_{P}=\left[\begin{array}[]{l}\partial_{P}\Psi_{0}\\ \partial_{P}\Psi_{0}^{*}\end{array}\right].$ (105)
We have explicitly evaluated the extended Berry-Gibbs curvature $F_{ab}=i(z_{a}|z_{b})$, recovering the predicted symplectic form, $F_{ab}=-\Omega_{ab}$. Specifically, in our gauge choice, the Berry-Gibbs curvature automatically vanishes without the need of further generalized gauge transformations, $F_{NP}=F_{PN}=0$ [see Eq. (191) and ensuing discussion]. In this example, a generalized gauge transformation amounts to add global phase and spatial shifts that depend on the particle number and momentum, $\theta=\theta(N,P),~{}x_{0}=x_{0}(N,P)$. The quantum amplitudes of the NG modes can be then understood as the quantum fluctuations of the global phase and position of the cnoidal wave, $\hat{X}^{\theta}=\delta\hat{\theta}$, $\hat{X}^{x}=\delta\hat{x}_{0}$, while their momenta describe total particle number and momentum fluctuations, $\hat{P}^{N}=\delta\hat{N}$, $\hat{P}^{P}=\delta\hat{P}$.
Regarding thermodynamics, in all numerical cases we have found that the matrix $\partial^{2}_{AB}E$ is both non-diagonal and not positive definite. This reflects, respectively, that both phase and crystalline modes hybridize in the low-frequency limit, and the presence of a negative-energy branch in the regular Bogoliubov sector as the true energy minimum is always an homogeneous GP solution Martone et al. (2021). In fact, the results of Ref. Martone et al. (2021), derived within the standard mean-field GP approach, correspond to setting $v=0$ instead of fixing the total momentum $P$, and taking the limit $L\to\infty$. However, even for $v=0$, the full GPG equation is still needed for the computation of $\partial^{2}_{AB}E$ and the subsequent characterization of the NG modes [see discussion after Eq. (81)].
@@
with a period $T=a/v$ and a quasi-chemical potential $\tilde{\mu}=\mu+vw_{\mathbf{n}}+v^{2}$, defined modulo $\omega=2\pi/T$. However, this time-periodic behavior does not stem from a genuine spontaneous symmetry breaking of time-translational invariance but rather from that of spatial-translational invariance plus a constant motion along that closed broken-symmetry orbit, as given by a non-vanishing $v$, in the same fashion of Eq. (24). Indeed, the NG mode arising from the alleged time-translation symmetry breaking is not independent from those of phase and spatial translations since $\partial_{t}\Psi=-i\mu\Psi-v\partial_{x}\Psi$, which implies $z_{t}=\mu z_{\theta}+vz_{x}$. This is also reflected by the energy dependence on the other conserved charges, $E=E(N,P)$. In general, for stationary GPG solutions, time translation symmetry is not spontaneously broken by itself but rather in terms of other symmetries as $\partial_{t}\Psi=-i(\boldsymbol{\lambda}\cdot\mathbf{T})\Psi$ and $E=E(\mathbf{Q})$.
This analysis suggests that certain time-periodic systems that do not genuinely break time-translational symmetry can be misidentified as continuous time crystals, with the time periodicity simply arising from constant motion along a closed orbit spawned by other broken generators.
V Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
In order to study spontaneous symmetry breaking of time-translation symmetry, we first analyze Floquet systems, which are described by periodic Hamiltonians explicitly breaking this symmetry, $\hat{H}(t+T)=\hat{H}(t)$. The periodic driving may not only be applied in the external potential Lignier et al. (2007); Kierig et al. (2008), but also in the interacting Gaul et al. (2011); Rapp et al. (2012) and even in the kinetic term Stöferle et al. (2004). In a many-body system whose dynamics can be described by a variational ansatz of the form (4), a periodic Hamiltonian is translated into
$i\frac{dX^{i}}{dt}=H^{i}_{j}(X,t)X^{j},$ (108)
where the matrix $H^{i}_{j}(X,t)$ satisfies $H^{i}_{j}(X,t+T)=H^{i}_{j}(X,t)$. We assume that the problem admits Floquet solutions of the form
$X^{i}(t)=\bar{X}^{i}(t)e^{-i\epsilon^{i}t},~{}\bar{X}^{i}(t+T)=\bar{X}^{i}(t),$ (109)
@@ V Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
$\epsilon^{i}\bar{X}^{i}(t)=\left[H^{i}_{j}(\bar{X}(t),t)-i\frac{d}{dt}\delta^{i}_{j}\right]\bar{X}^{j}(t),$ (110)
where $H^{i}_{j}(\bar{X}(t),t)$ is now a fully periodic operator as the vector $\bar{X}(t)$ is also periodic. The dynamics of the collective modes is described by a linear equation analogous to Eq. (6),
$i\frac{d\delta\bar{X}^{i}}{dt}=M^{i}_{j}(t)\delta\bar{X}^{j},$ (111)
with $M^{i}_{j}(t+T)=M^{i}_{j}(t)$. As a linear periodic equation, it admits Floquet solutions of the form $\delta\bar{X}^{i}(t)=u^{i}_{\varepsilon}(t)e^{-i\varepsilon t}$, with $u^{i}_{\varepsilon}(t+T)=u^{i}_{\varepsilon}(t)$ and $\varepsilon$ the quasi-energy. The corresponding eigenvalue equation reads
@@
$0=\left[M^{i}_{j}(t)-i\frac{d}{dt}\delta^{i}_{j}\right](L\bar{X})^{j},$ (113)
so $L\bar{X}$ is a Floquet mode with zero quasi-energy. We can identify it as the NG mode associated to the spontaneous-symmetry breaking, and consequently we will denote it as Floquet-Nambu-Goldstone mode.
It was recently shown de Nova and Sols (2022) that, due to many-body interactions, Floquet states can spontaneously emerge even in the absence of external periodic driving as self-consistent Floquet solutions (109) to the original time-independent problem (4),
$\epsilon^{i}\bar{X}^{i}(t)=\left[H^{i}_{j}(\bar{X}(t))-i\frac{d}{dt}\delta^{i}_{j}\right]\bar{X}^{j}(t).$ (114)
It is important to remark that here the periodicity of $H^{i}_{j}(\bar{X}(t))$ stems solely from that of $\bar{X}(t)$, and is spontaneously determined by the solution itself and not externally imposed, in contrast to Eq. (110). The linear problem is also described by Floquet modes, as in Eq. (112), but once more the periodicity of the matrix $M^{i}_{j}(t)$ stems just from that of $\bar{X}(t)$. Similarly, spontaneous symmetry breaking is translated into the emergence of FNG modes. However, a spontaneous Floquet state also breaks time-translation symmetry as $\bar{X}(t+t_{0})$ is as well a solution of Eq. (114). This implies that $\partial_{t}\bar{X}(t)$ is a FNG mode,
$0=\left[M^{i}_{j}(t)-i\frac{d}{dt}\delta^{i}_{j}\right]\partial_{t}\bar{X}^{j}(t),$ (115)
@@
VI Simultaneous symmetry breaking in spontaneous Floquet states
We directly adapt the formalism developed in Sec. III to study simultaneous symmetry breaking in a spontaneous Floquet state, focusing once more on the case of a condensate for simplicity. Within this framework, spontaneous Floquet states are periodic solutions

### v4

@@ 4 Cnoidal wave in a superfluid
As a particular application of the formalism developed in the previous section, we consider a 1D quasi-condensate [81] in a ring of length $L$ with no external potential, $V(x)=0$, so there is invariance under spatial translations. This type of model has been recently used to understand certain aspects of supersolidity due to its appealing simplicity [57]. We now look for solutions which simultaneously break both $U(1)$ and space-translation symmetry. In our notation, this amounts to taking $\boldsymbol{\alpha}=(\theta,x)$, $\mathbf{T}=(1,-i\partial_{x})$ and $\mathbf{Q}=(N,P)$, respectively. The Lagrange multipliers associated to the particle number $N$ and momentum $P$ are $\lambda^{\theta}=\mu$ and $\lambda^{x}=v$, i.e., $\boldsymbol{\lambda}=(\mu,v)$. The resulting GPG equation (22) then reads
@@
$H_{GP}\tilde{\Psi}_{0}=\left(\mu+\frac{v^{2}}{2}\right)\tilde{\Psi}_{0}\equiv\mu_{v}\tilde{\Psi}_{0}.$ (88)
@@
$\theta(x)$ $=$ $\int^{x}_{0}\frac{J}{n(x^{\prime})}\mathrm{d}x^{\prime}=\frac{J}{n_{1}\sqrt{n_{3}-n_{1}}}\Pi\left(\text{am}(\sqrt{n_{3}-n_{1}}x,\nu),m,\nu\right),$
@@
$\nu=\frac{n_{2}-n_{1}}{n_{3}-n_{1}},\leavevmode\nobreak\ m=1-\frac{n_{2}}{n_{1}},$ (90)
@@
$\mu_{v}=\frac{n_{1}+n_{2}+n_{3}}{2},\leavevmode\nobreak\ J=\sqrt{n_{1}n_{2}n_{3}}.$ (91)
@@
Cnoidal waves are periodic solutions of a nonlinear wave equation, generalizing the usual linear sinusoidal waves. The simplest example is the oscillation of a pendulum, which can be analytically obtained in terms of elliptic functions [see Eq. (A.1) and ensuing discussion]. In our case, for fixed $J,\mu_{v}$, the parameter $\nu$ controls both the periodicity and the nonlinearity of the wave; see Fig. 1. In the linear limit $\nu\to 0$, we recover a small amplitude sinusoidal (horizontal black and solid blue lines), $\textrm{sn}(u,0)=\sin u$ and $K(0)=\pi/2$, as in the case of the small oscillations of a pendulum around its stable equilibrium position. For increasing $\nu$, the oscillation amplitude and periodicity grow (dashed-dotted green and dashed red); in the limit $\nu\to 1$, the period diverges as $\textrm{sn}(u,1)=\tanh u$ and $K(1)=\infty$. This is akin to the falling trajectory of a pendulum whose starting point is close to the unstable equilibrium position.
Specifically, the period $a$ of the cnoidal density $n(x)$ is
@@
Regarding the phase $\theta(x)$, since $n(x)$ is periodic, $J/n(x)$ is periodic. The integral of a periodic function is a periodic function modulo a constant slope $w_{\mathbf{n}}x$, given by its zero Fourier component. By taking $\theta(0)=0$, we have
@@
$L$ $=$ $\ell a=\frac{2K(\nu)\ell}{\sqrt{n_{3}-n_{1}}},\leavevmode\nobreak\ \ell\in\mathbb{N},$ (96)
$\frac{2\pi q}{L}$ $=$ $v+w_{\mathbf{n}},\leavevmode\nobreak\ q\in\mathbb{Z},$
$N$ $=$ $\int^{L}_{0}\mathrm{d}x\leavevmode\nobreak\ |\Psi_{0}(x)|^{2},$
$P$ $=$ $\int^{L}_{0}\mathrm{d}x\leavevmode\nobreak\ \Psi^{*}_{0}(-i\partial_{x})\Psi_{0},$
@@
$\frac{2\pi q}{L}$ $=$ $\frac{\bar{p}}{\bar{n}}+\sqrt{n_{1}n_{2}n_{3}}\left[\frac{\Pi(m,\nu)}{n_{1}K(\nu)}-\frac{1}{\bar{n}}\right].$
@@
$\bar{n}\equiv\frac{N}{L},\leavevmode\nobreak\ \bar{p}\equiv\dfrac{P}{L}.$ (98)
The procedure to numerically solve these equations is explained after Eq. (205). Once done, we evaluate the energy density by
@@
and numerically compute the derivatives $\partial_{N}E=\partial_{\bar{n}}e$, $\partial_{P}E=\partial_{\bar{p}}e$, checking that indeed we recover $\partial_{N}E=\mu,\leavevmode\nobreak\ \partial_{P}E=v$. We note that the above formalism can be straightforwardly adapted to account for the presence of a flux within the ring.
@@
$z_{\theta}(x)$ $=\left[\begin{array}[]{r}-i\Psi_{0}(x)\\ i\Psi_{0}^{*}(x)\end{array}\right],\leavevmode\nobreak\ z_{x}(x)=-\left[\begin{array}[]{l}\partial_{x}\Psi_{0}(x)\\ \partial_{x}\Psi_{0}^{*}(x)\end{array}\right],$ (104)
$z_{N}(x)$ $=\left[\begin{array}[]{l}\partial_{N}\Psi_{0}(x)\\ \partial_{N}\Psi_{0}^{*}(x)\end{array}\right],\,z_{P}(x)=\left[\begin{array}[]{l}\partial_{P}\Psi_{0}(x)\\ \partial_{P}\Psi_{0}^{*}(x)\end{array}\right].$ (109)
We have explicitly evaluated the extended Berry-Gibbs curvature $F_{ab}=i(z_{a}|z_{b})$, recovering the predicted symplectic form, $F_{ab}=-\Omega_{ab}$. Specifically, in our gauge choice, the Berry-Gibbs curvature automatically vanishes without the need of further generalized gauge transformations, $F_{NP}=F_{PN}=0$ [see Eq. (206) and ensuing discussion]. In this example, a generalized gauge transformation amounts to adding global phase and spatial shifts that depend on the particle number and momentum, $\theta=\theta(N,P),\leavevmode\nobreak\ x_{0}=x_{0}(N,P)$. The quantum amplitudes of the NG modes can be then understood as the quantum fluctuations of the global phase and position of the cnoidal wave, $\hat{X}^{\theta}=\delta\hat{\theta}$, $\hat{X}^{x}=\delta\hat{x}_{0}$, while their momenta describe total particle number and momentum fluctuations, $\hat{P}^{N}=\delta\hat{N}$, $\hat{P}^{P}=\delta\hat{P}$.
Regarding thermodynamics, in all numerical cases we have found that the susceptibility matrix $\partial^{2}_{AB}E$ is both non-diagonal and non-positive definite. This reflects, respectively, that phase and crystalline modes hybridize in the low-energy limit, and the presence of a negative-energy branch in the regular Bogoliubov sector as the true energy minimum is always a homogeneous GP solution [57]. In fact, the results of Ref. [57], derived within the standard mean-field GP approach, correspond to setting $v=0$ instead of fixing the total momentum $P$, and taking the limit $L\to\infty$. However, even for $v=0$, the full GPG equation is still needed for the computation of the susceptibility matrix and the subsequent characterization of the Goldstone-Gibbs modes [see discussion after Eq. (85)].
@@
with a period $T=a/v$ and a quasi-chemical potential $\tilde{\mu}=\mu+vw_{\mathbf{n}}+v^{2}$, defined modulo $\omega=2\pi/T$. However, this time-periodic behavior does not stem from a genuine spontaneous symmetry breaking of time-translation invariance but rather from that of space-translation invariance plus a constant motion along a closed broken-symmetry orbit, as set by a non-vanishing $v$, in the same fashion of Eq. (23). Indeed, the NG mode arising from the alleged time-translation symmetry breaking is not independent from those of phase and spatial translations since $\partial_{t}\Psi=-i\mu\Psi-v\partial_{x}\Psi$, which implies $z_{t}=\mu z_{\theta}+vz_{x}$. This is also reflected by the energy dependence on the other conserved charges, $E=E(N,P)$. In general, for stationary GPG solutions, time-translation symmetry is not spontaneously broken by itself but rather in terms of other symmetries as $\partial_{t}\Psi=-i(\boldsymbol{\lambda}\cdot\mathbf{T})\Psi$ and $E=E(\mathbf{Q})$.
This analysis suggests that certain time-periodic systems that do not genuinely break time-translation symmetry can be misidentified as continuous time crystals, with the time periodicity simply arising from constant motion along a closed orbit spawned by other broken-symmetry generators.
5 Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
In order to study spontaneous symmetry breaking of time-translation symmetry, we first analyze Floquet systems, which are described by periodic Hamiltonians explicitly breaking this symmetry, $\hat{H}(t+T)=\hat{H}(t)$. The periodic driving may not only be applied in the external potential [82, 83], but also in the interacting [84, 85] and even in the kinetic term [86, 41]. In a many-body system whose dynamics can be described by a variational ansatz of the form (4), a periodic Hamiltonian is translated into
$i\frac{dX^{i}}{dt}=H^{i}_{j}(\mathbf{X},t)X^{j},$ (113)
where the matrix $H^{i}_{j}(\mathbf{X},t)$ satisfies $H^{i}_{j}(\mathbf{X},t+T)=H^{i}_{j}(\mathbf{X},t)$. We assume that the problem admits Floquet solutions of the form
$X^{i}(t)=X_{0}^{i}(t)e^{-i\epsilon^{i}t},\leavevmode\nobreak\ X_{0}^{i}(t+T)=X_{0}^{i}(t),$ (114)
@@ 5 Variational Floquet-Goldstone Theorem: Floquet-Nambu-Goldstone modes
$\epsilon^{i}X_{0}^{i}(t)=\left[H^{i}_{j}(\mathbf{X}_{0}(t),t)-i\frac{d}{dt}\delta^{i}_{j}\right]X_{0}^{j}(t),$ (115)
where $H^{i}_{j}(\mathbf{X}_{0}(t),t)$ is now a fully periodic operator as the vector $\mathbf{X}_{0}(t)$ is also periodic. The dynamics of the collective modes is described by a linear equation analogous to Eq. (6),
$i\frac{d\delta X^{i}}{dt}=M^{i}_{j}(t)\delta X^{j},$ (116)
with $M^{i}_{j}(t+T)=M^{i}_{j}(t)$. As a linear periodic equation, it admits Floquet solutions of the form $\delta X^{i}(t)=u^{i}_{\varepsilon}(t)e^{-i\varepsilon t}$, with $u^{i}_{\varepsilon}(t+T)=u^{i}_{\varepsilon}(t)$ and $\varepsilon$ the quasi-energy. The corresponding eigenvalue equation reads
@@
$0=\left[M^{i}_{j}(t)-i\frac{d}{dt}\delta^{i}_{j}\right]Y^{j},$ (118)
so $\mathbf{Y}=-iL\mathbf{X}_{0}$ is a Floquet mode with zero quasi-energy. We can identify it as the NG mode associated to the spontaneous-symmetry breaking, and consequently we will denote it as Floquet-Nambu-Goldstone mode.
It was recently shown [43] that, due to many-body interactions, Floquet states can spontaneously emerge even in the absence of external periodic driving as self-consistent Floquet solutions (114) to the original time-independent problem (4),
$\epsilon^{i}X_{0}^{i}(t)=\left[H^{i}_{j}(\mathbf{X}_{0}(t))-i\frac{d}{dt}\delta^{i}_{j}\right]X_{0}^{j}(t).$ (119)
It is important to remark that here the periodicity of $H^{i}_{j}(\mathbf{X}_{0}(t))$ stems solely from that of $\mathbf{X}_{0}(t)$, and is spontaneously determined by the solution itself and not externally imposed, in contrast to Eq. (115). The linear problem is also described by Floquet modes, as in Eq. (117), but once more the periodicity of the matrix $M^{i}_{j}(t)$ stems just from that of $\mathbf{X}_{0}(t)$. Similarly, spontaneous symmetry breaking is translated into the emergence of FNG modes. However, a spontaneous Floquet state also breaks time-translation symmetry since $\mathbf{X}_{0}(t+t_{0})$ is as well a solution of Eq. (119). This implies that $\mathbf{Y}=\partial_{t}\mathbf{X}_{0}$ is an FNG mode,
$0=\left[M^{i}_{j}(t)-i\frac{d}{dt}\delta^{i}_{j}\right]\partial_{t}X_{0}^{j},$ (120)
@@
6 Simultaneous symmetry breaking in spontaneous Floquet states
We directly adapt the formalism developed in Sec. 3 to study simultaneous symmetry breaking in spontaneous Floquet states, focusing once more on the case of a condensate for simplicity. Within this framework, spontaneous Floquet states are periodic solutions